Consider transverse waves moving along a stretched spring Suppose the mass of the spring is 10 grams, its relaxed (not stretched) length is 1.00 meters, and its spring constant (if stretched) is 5.0 N/cm. The spring is tensioned via the pulley wheel and a hanging weight (as in the experiment). The final length of the stretched spring is 1.50 meters (from the oscillator to the weight). According to Eqn 2, how fast will transverse waves propagate along this stretched spring? SHOW YOUR WORK.
Answer #1


Tension in the spring is F = k*x

F = 500 N/m*(1.5-1) = 500*0.5 = 250 N

linear density is mu= m/l = 0.010/1 = 0.01 kg/m

speed of the transverse wave is v =Sqrt(F/mu) = Sqrt(250/0.01) = 158.11 m/s